# Towards A Universal Graph Structural Encoder

### v1

@@ 4.4.2 Evaluation on Molecular Graphs
Table 3: Test ROC-AUC(%) performance on molecular datasets. The best results for each dataset are bolded.
@@
4.5 Integration with Large Language Models
Settings and Methods. We perform experiments on citation networks and e-commerce networks, including Arxiv (Hu et al., 2020a), Cora (Yang et al., 2016) and Amazon-product (He & McAuley, 2016; McAuley et al., 2015), which are text-attributed graphs with detailed descriptions for each node (e.g., product item). Edges indicate co-authored or co-purchased relations between two nodes. We employ a lightweight MLP to align the PSE generated by GFSE with the language model’s embedding space, which ensures seamless integration of structural information into the language model. We concatenate the textual description of a central node with those of its one-hop neighbors and prepend the PSE as a soft token, followed by a special graph token at the end. This combined sequence is then encoded by a language model. The hidden embedding of the special graph token is used as the representation for the central node. Following the previous setting (Zhu et al., 2024), we compute the cosine similarity between the representations of node pairs as the edge likelihood. We train the MLP and fine-tune the language model with LoRA (Hu et al., 2021) using a contrastive loss (Hadsell et al., 2006). More evaluation details can be found in Appendix E.2.
Table 4: Evaluation on text-attributed graphs
@@ 4.5 Integration with Large Language Models
InstructGLM 76.23 82.60 79.82 85.93 62.50 73.25 74.94 79.95
Finetuned LLaMA 74.73 82.87 78.93 86.07 62.52 75.77 75.70 87.08
@@
Results. Evaluation results on text-attributed graphs are reported in Table 4. We select InstructGLM (Ye et al., 2023) as a baseline, which has been fine-tuned on graph domains without structural information infusion. Additionally, we include comparisons against RRWP (Ma et al., 2023), a random-walk based PSE, and GraphSAGE (Hamilton et al., 2017), which was trained from scratch as a PSE encoder with LLM finetuning together. Finetuned LLaMA refers to the LLaMA model (Touvron et al., 2023) fine-tuned without any PSE. As seen from the table, GFSE, which is pre-trained on cross-domain graph data as a structural encoder, outperforms other methods in general. In particular, GFSE achieves an average performance boost of $2.66\%$ over InstructGLM, $2.02\%$ over GraphSAGE encoder and $1.53$ over RRWP across the five datasets. These gains demonstrate the effectiveness of GFSE in language-based graph tasks. See more discussions in Appendix F.1.
4.6 More Experimental Analysis
Table 5: Ablation studies on the pre-training tasks and model architecture of GFSE. Best results are shown in bold.
@@ 4.6 More Experimental Analysis
Ablation on Tasks and Model Architecture. We analyze the sensitivity of each pre-training task and model architecture w.r.t. the performance boost in downstream tasks. The results are illustrated in Table 4.6. Firstly, we iteratively remove one pre-training task and follow the same setting to pre-train and evaluate GFSE. We observe that each removal results in a discernible reduction in downstream performance. Notably, the shortest path distance task is particularly critical for the ZINC and CIFAR10 datasets, while community detection is significant for academic datasets.
We further conduct ablation studies on the main components of GFSE. Specifically, we use traditional attention to replace biased attention as a baseline and remove GIN or attention modules respectively. We notice that all the above pre-training architecture variants lead to performance degradation. The hybrid approach outperforms both attention-only and GIN-only setups, suggesting that integrating sophisticated attention mechanisms can compensate for the absence of global information in local message-passing layers.
@@
Task-specific Uncertainty. Figure 3(a) illustrates the trajectory of task uncertainty ($\sigma^{2}$) across different pre-training tasks w.r.t. pre-training epochs. Higher values of $\sigma^{2}$ reduce the respective task’s contribution to the overall training loss. We observe that all tasks show a sharp decline in uncertainty during the pre-training process. Notably, motif counting maintains a lower uncertainty throughout the training process compared to other tasks, suggesting that it might be inherently more straightforward for the model to optimize or more integral to the model’s overall learning structure.
Figure 3: (a) Learning task uncertainty ($\sigma^{2}$) w.r.t. pre-training. (b) Performance of GFSE pre-trained on different data sizes. Evaluation datasets are MolPCBA and Arxiv. The base model is GPS.
@@
Impact of Pre-training Data Size. Figure 3(b) illustrates the GFSE’s performance when pre-trained on varying ratios of data, revealing several key insights. (1) The model performance declines as the pre-training data size decreases, highlighting the importance of a diverse and extensive pre-training dataset for enhancing the model effectiveness. (2) Pre-training is a critical factor for achieving superior and expressive graph representations, as evidenced by the consistent outperformance of pre-trained GFSE over its train-from-scratch counterpart. (3) Cross-domain pre-training proves to be important for domain generalization. GFSE pre-trained on a single dataset struggles to generalize to different domains (e.g., from MolPCBA to Arxiv). This indicates that models trained on a narrow domain may overfit to the specific characteristics, limiting their generalization to diverse downstream tasks. In contrast, cross-domain pre-training enables the model to learn more transferable and robust representations, which are essential for achieving strong performance across a wide range of applications.
5 Conclusion
GFSE is a significant achievement in universal graph encoder, leveraging multiple self-supervised learning objectives to capture comprehensive structural information from diverse graph domains. By integrating relative positional encoding within a Graph Transformer, GFSE provides a robust framework for generating expressive positional and structural encodings. Extensive experiments on synthetic and real-world datasets validate GFSE’s ability to enhance the performance of various graph feature encoders, broadening its applicability across numerous graph-related tasks. Building upon the promising results of GFSE, a promising future direction could involve dynamic graph structures in the model pre-training, allowing it to adapt to evolving graph topologies over time.
Appendix A Positional and Structural Encoding
Random Walk Structural Encoding. Random walk structural encoding (RWSE) (Dwivedi et al., 2021; Rampášek et al., 2022) uses landing probabilities derived from random walks of varying lengths, starting from each node in the graph to capture both structural and positional relationships among nodes. Formally, let $G(V,E)$ represent an input graph, where $V$ and $E$ denote the set of $N$ nodes and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ indicates the adjacency matrix and $\mathbf{D}$ is the degree matrix. Random Walk matrix is defined as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{M}_{i,j}$ indicates the transition probability from the $i$-th node to the $j$-th node. The RWSE with $d$ steps for the $i$-th node is defined as
$\text{RWSE}_{i}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d}]_{i,i},$ (6)
Laplacian Positional Encoding. Laplacian positional encoding (LapPE) (Kreuzer et al., 2021; Dwivedi et al., 2023) method emerges as a significant advancement for enriching node representations with spectral information. LapPE utilizes the eigenvectors of the graph Laplacian matrix to encode the relative positions of nodes within a graph. These eigenvectors reflect a local coordinate system that captures meaningful structural information, while also preserving the global topological properties of the original graph. The Laplacian matrix $\mathbf{L}=\mathbf{D}-\mathbf{A}$ has the full eigendecomposition as $\mathbf{L}=\mathbf{U\Lambda U}^{\top}$. We use the $k$ smallest non-trivial eigenvectors of the Laplacian matrix to create the LapPE. The LapPE with $k$ eigenvectors for the $i$-th node is expressed as:
$\text{LapPE}_{i}=\left[\boldsymbol{u}_{1,i},\boldsymbol{u}_{2,i},\cdots,\boldsymbol{u}_{k,i}\right]\in\mathbb{R}^{k},$ (7)
where $\boldsymbol{u}_{t}$ denotes the $t$-th smallest non-trivial eigenvectors and $k$is the number of eigenvectors used. Laplacian PE is particularly useful in situations where nodes are inherently anonymous and lack unique features. However, the Laplacian encoding faces challenges from the arbitrary sign of normalized eigenvectors, introducing $2^{k}$ possible configurations for $k$ eigenvectors. To manage this complexity during training, eigenvectors are randomly sampled from these possibilities. Alternatively, resolving the sign ambiguity by taking the absolute values of eigenvectors simplifies the model but can significantly reduce the expressiveness of the positional features.
Pre-trained Positional and Structural Encoder. Graph Positional and Structural Encoder (GPSE) (Liu et al., 2023b) is a graph encoder pre-trained on molecule datasets by reconstructing traditional positional encoding, such as LapPE, RWSE, CycleSE, etc. The model takes as input graph adjacency matrix and randomly generates node features by $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to improve the expressiveness. The pre-training architecture is a deep MPNN with 20 layers and residual connection and gating mechanism. Moreover, GPSE utilizes a virtual node technique in each graph to enable global message passing. However, GPSE suffers from poor generalizability across other domains, due to its pretraining setting and randomized node features.
Appendix B Dataset Pre-processing
B.1 Shortest Path Distance Regression
Shortest Path Distance (SPD) Regression is an edge-level reconstruction task. SPD encodes the global proximity and connectivity between nodes, which helps to discern nodes’ positions and relations within the entire graph (Li et al., 2020). For data preprocessing, we pre-calculate the $N\times N$ SPD matrix of a given graph before the pretraining phase and save the matrix with the graph data for fast retrieval. We utilize the Dijkstra algorithm (Dijkstra, 1959) to compute the shortest path distances between node pairs, which serve as SPD labels. During pretraining, we randomly select node pairs and fetch their SPD as labels to perform the edge-level reconstruction task. The loss for shortest path distance regression is computed as:
$\mathcal{L}_{\text{SPD}}=\frac{1}{|E|}\sum_{i,j\in V}\|h_{\text{SPD}}(\mathbf{P}^{L}_{i}\|\mathbf{P}^{L}_{j})-\text{SPD}_{i,j}\|^{2}$
[…]
B.2 Motif Counting
Motif Counting is formulated as a node-level reconstruction task. We explicitly use subgraph isomorphism counting as a form of self-supervision during pretraining, allowing the model to better leverage structural information and identify each node’s role in the surrounding subgraphs (Bouritsas et al., 2022). We follow previous works on subgraph isomorphism to count the number of certain motifs surrounding each node. A key concept in our preprocessing step is ”Automorphism orbits,” as introduced by Pržulj (Pržulj, 2007). This concept helps in identifying unique roles of nodes within the counted motifs. The detailed counting method is as follows:
  1. 1.
    Define the subgraph structures to be counted: Enumerate all graphlets with the number of nodes less than or equal to 4. This limitation is due to the exponential growth in the number of graphlets with increasing nodes, which becomes computationally infeasible.
  2. 2.
    Assign indexed orbits: For each subgraph (or motif), assign each vertex a uniquely indexed orbit to facilitate accurate counting.
  3. 3.
    Count specified subgraphs: Utilize the subgraph_isomorphism function from the Python package graph-tool to count specified subgraphs throughout the entire graph. This count is used to determine orbits for each node in the entire graph, forming an orbit degree vector.
  4. 4.
    Save and prepare node-level labels: Save these vectors with the data as node-level labels, which are then prepared for our node-level reconstruction task.
Figure 4: Illustration of various subgraphs (graphlets) used in the motif counting. Each subgraph is indexed and labeled for reference.
Appendix C Methodologies
C.1 Dataset
@@
Table 6: Dataset Information. class. represents classification task and reg. represents regression task.
@@
MolPCBA dataset consists of 437,929 molecular graphs, each representing a compound from the PubChem BioAssay database. This dataset is used for multi-label binary classification tasks across 128 targets, focusing on predicting the bioactivity of compounds against various protein targets. The primary evaluation metric for this dataset is Average Precision (AP).
MolHIV contains 41,127 molecular graphs derived from the MoleculeNet benchmark. Each graph represents a molecule, with nodes as atoms and edges as chemical bonds. The task is a binary classification to predict the ability of molecules to inhibit HIV replication, with AUROC as the evaluation metric.
MNIST dataset includes 70,000 images converted into graphs. Each image represents a handwritten digit, with nodes representing pixels and edges representing pixel adjacency. The task is a 10-way classification to identify the digit in the image, evaluated using accuracy (ACC).
Peptides-func comprises 15,535 graphs, each representing a peptide. Nodes represent amino acids, and edges represent peptide bonds. The task involves binary classification to predict the functional properties of the peptides, with Average Precision (AP) as the evaluation metric.
Peptides-struct also contains 15,535 peptide graphs but focuses on regression tasks to predict structural properties of peptides, such as bond angles and distances. The evaluation metric is Mean Absolute Error (MAE).
ogbn-proteins dataset is a large-scale graph with 132,534 nodes and 39,561,252 edges, representing protein-protein interaction networks. Each node is a protein, and edges represent interactions. The task is binary classification at the node level to predict protein functions, evaluated using AUROC.
Pokec is a social network dataset from the Pokec online social network in Slovakia. It includes one large graph with 1,632,803 nodes (users) and 30,622,564 edges (friendships). The task is binary classification to predict user attributes, such as gender, with accuracy (ACC) as the metric.
ogbn-arxiv dataset consists of a single large graph with 169,343 nodes and 1,166,243 edges, representing the citation network of arXiv papers. Each node is a paper, and edges represent citation links. The task is 40-way classification to predict the primary subject area of each paper, evaluated using accuracy (ACC).
ogbn-products dataset includes a large graph with 2,449,029 nodes and 61,859,140 edges, representing an Amazon product co-purchasing network. Nodes represent products, and edges represent co-purchasing relationships. The task is 47-way classification to predict the product category, with accuracy (ACC) as the metric.
ZINC contains 249,456 molecular graphs, where each graph represents a molecule from the ZINC database. The task is regression to predict molecular properties like solubility, with Mean Absolute Error (MAE) as the evaluation metric.
PubMed dataset includes a citation network of 19,717 scientific publications. Nodes represent papers, and edges represent citations. The task is 3-way classification to predict the subject areas of the papers, evaluated using accuracy (ACC).
CIFAR10 dataset has 60,000 images transformed into graphs, where each image represents a colored object. Nodes represent pixels, and edges represent pixel adjacency. The task is a 10-way classification to identify the object in the image, evaluated using accuracy (ACC).
Triangle poses triangle counting as a 10-way graph-level classification task. Half of the test set are graphs with a similar size to those in the training and validation set (denoted as Triangle-S). The left are graphs with larger sizes (denoted as Triangle-L), which present greater challenges to the model’s expressive power.
Pattern and Cluster are graph datasets generated with the Stochastic Block Model (SBM) (Abbe, 2018) that require the model to discern graph patterns and local clusters, which is a node-level classification task.
Table 7: Dataset statistics of three categories from Amazon e-commerce networks
@@ C.2 Pre-training Metric
  • For the community detection task, we set $\epsilon$ as $1$. The predicted label between the $i$-th node and the $j$-th node $\hat{Y}_{i,j}^{c}$ is $1$ if $\operatorname{sim}(i,j)\geq 0.5$ and 00 otherwise. Accuracy is calculated by comparing the predicted label $\hat{Y}_{i,j}^{c}$ with the ground truth labels $Y_{i,j}^{c}$ and is defined as the proportion of correctly predicted labels out of all possible node pairs:
[…]
    where $\mathbbm{1}(\cdot)$ is an indicator function and $|V|(|V|-1)/2$ is the total number of unique node pairs in the graph. This metric effectively measures how well the model can identify community structures by correctly classifying node pairs as being in the same community or in different communities.
@@
[…]
    where $\hat{Y}_{G_{i},G_{j}}$ is the predicted label indicating whether graph $G_{i}$ and $G_{j}$ are from the same dataset and $Y_{G_{i},G_{j}}$ is the ground truth label. $\hat{Y}_{G_{i},G_{j}}$ is $1$ if $\operatorname{sim}(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}})\geq 0$ and 00 otherwise.$N$ is the total number of evaluated graph pairs.
@@
[…]

### v2

@@ 4.4.2. Evaluation on Molecular Graphs
Table 4. Test ROC-AUC(%) performance on molecular datasets. The best results for each dataset are bolded.
@@
Table 5. Evaluation on text-attributed graphs
@@
InstructGLM 76.23 82.60 79.82 85.93 62.50 73.25 75.70 87.08
Finetuned LLaMA 74.73 82.87 78.93 86.07 62.52 75.77 74.94 79.95
@@
4.5. Integration with Large Language Models
We evaluate GFSE on the task of link prediction using three text-attributed graphs: the citation networks Arxiv (hu2020open) and Cora (yang2016revisiting), and the e-commerce network Amazon-Product (he2016ups; mcauley2015image). In these datasets, each node (e.g., a paper or product) features a detailed textual description, and edges signify relations such as co-authorship or co-purchase. To synergize our structural encodings with an LLM, we employ a lightweight MLP to project the 32-dimensional PSE into the 4096-dimensional embedding space of LLaMA2-7B (touvron2023llama), aligning the structural and textual modalities. The input for the LLM is constructed by concatenating the central node’s text with that of its one-hop neighbors. This sequence is then prepended with the projected PSE as a soft structural token and appended with a special graph token. The final hidden state corresponding to this special token serves as the comprehensive node representation. Following a prior setup (zhu2024parameter), we train the model to predict edge likelihoods based on the cosine similarity between node-pair representations. Specifically, we train the MLP projector and fine-tune the LLM using LoRA (hu2021lora) with the contrastive objective: $\mathcal{L}=\sum_{(i,j)\in E\&(i,j^{\prime})\notin E}\big(d_{ij}^{2}+\max(\tau-d_{ij^{\prime}},0)^{2}\big)$, where $d_{ij}=1-\cos(v_{i},v_{j})$, $v_{i}$ denotes the $i$-th node representation, and $\tau=0.5$ is the margin.
Results. Evaluation results on text-attributed graphs are reported in Table 5. We select InstructGLM (ye2023natural) as a baseline, which has been fine-tuned on graph domains without structural information. Additionally, we include comparisons against RRWP (ma2023graph), and GraphSAGE (hamilton2017inductive), which are trained from scratch as a PSE encoder with LLM finetuning together. Finetuned LLaMA refers to the LLaMA model (touvron2023llama) fine-tuned without any PSE. As seen from the table, GFSE, which is pre-trained on cross-domain graph data as a structural encoder, outperforms other methods in general. In particular, GFSE achieves an average performance boost of $2.66\%$ over InstructGLM, $2.02\%$ over GraphSAGE encoder and $1.53$ over RRWP across the five datasets. These gains demonstrate the effectiveness of GFSE in language-based graph tasks. See more discussions in Appendix B.2.
4.6. More Experimental Analysis
Reconstruction of other PSE types. The pre-trained GFSE is able to generate effective PSE, which can reconstruct other pre-defined PSE, such as LapPE (lim2022sign) and ElstaticPE (kreuzer2021rethinking). Table 6 demonstrates that the PSE generated by the pre-trained GFSE, followed by a trainable lightweight MLP, is capable of reconstructing various other pre-defined PSEs. We evaluate this using the coefficient of determination $R^{2}$ scores as a metric. Notably, our model performs competitively compared with GPSE, given the fact that GPSE directly adopts PSE reconstruction as its training objective. Instead, our method generalizes well across different PSEs without being directly trained for reconstruction. This suggests that the structural self-supervision pre-training tasks are effective and sufficient in capturing important structural information.
Table 6. Performance of other PSE reconstruction. The coefficient of determination $R^{2}$ scores are reported as the metric.
PSE type ElstaticPE LapPE RWSE HKdiagSE CycleSE
GPSE 0.964 0.973 0.984 0.981 0.977
Ours 0.947 0.986 0.987 0.984 0.992
Ablation on Tasks and Model Architecture. We analyze the sensitivity of each pre-training task and model architecture w.r.t. the performance boost in downstream tasks. The results are illustrated in Table 7. Firstly, we iteratively remove one pre-training task and follow the same setting to pre-train and evaluate GFSE. We observe that each removal results in a discernible reduction in downstream performance. Notably, the shortest path distance task is particularly critical for the ZINC and CIFAR10 datasets, while community detection is significant for academic datasets. We also conduct ablation studies on the main components of GFSE. Specifically, we use traditional attention to replace biased attention and remove GIN or attention modules, respectively. We notice that all the above architecture variants lead to performance degradation. The hybrid approach outperforms both attention-only and GIN-only setups, suggesting that integrating biased attention mechanisms can compensate for the absence of global information in local message-passing layers.
Table 7. Ablation studies on the pre-training tasks and model architecture of GFSE. Best results are shown in bold.
@@ 4.6. More Experimental Analysis
Figure 4. Learning task uncertainty ($\sigma^{2}$) w.r.t. pre-training
@@
Task-specific Uncertainty. Figure 4 illustrates the trajectory of task uncertainty ($\sigma^{2}$) across different pre-training tasks w.r.t. pre-training epochs. Higher values of $\sigma^{2}$ reduce the respective task’s contribution to the overall training loss. We observe that all tasks show a sharp decline in uncertainty during the pre-training process. Notably, motif counting maintains a lower uncertainty throughout the training process compared to other tasks, suggesting that it might be inherently more straightforward for the model to optimize or more integral to the model’s overall learning structure.
@@
Impact of Pre-training Data Size. Figure 5 illustrates the impact of pre-training data scale and domain transfer on GFSE’s performance, revealing several key insights. First, as the pre-training data ratio increases, the model’s performance consistently improves, indicating that larger and more diverse pre-training datasets are essential for capturing richer structural representations and achieving stronger generalization. Second, models initialized with GFSE pre-training outperform their train-from-scratch counterparts, demonstrating the effectiveness of structural pre-training in providing transferable inductive biases that enhance downstream adaptation. Lastly, cross-domain pre-training proves to be important for domain generalization. GFSE pre-trained on a single dataset struggles to generalize to different domains (e.g., from MolPCBA to Arxiv). This indicates that models trained on a single domain may overfit to the specific characteristics, limiting their generalization to other tasks. In contrast, cross-domain pre-training enables the model to learn more robust, domain-invariant representations, crucial for scaling to diverse real-world text-attributed graph applications.
Figure 5. Scalability of GFSE with GPS as a backbone.
5. Conclusion
GFSE represents a meaningful step toward a universal graph encoder, leveraging multiple self-supervised objectives and relative positional encoding within a Graph Transformer to capture expressive and transferable structural patterns across diverse graph domains. Extensive experiments on both synthetic and real-world datasets demonstrate that GFSE consistently enhances the performance of downstream graph encoders across a wide range of tasks. The effectiveness of GFSE can be influenced by the quality and diversity of the pre-training graphs, as biases or under-representation in these data sources may affect the learned representations. Incorporating richer, more heterogeneous, and potentially dynamic graph data is a natural direction for future research. More broadly, GFSE holds promise for real-world applications such as computational biology, social network analysis, and recommendation systems, while reducing reliance on heavy task-specific fine-tuning and improving the accessibility of strong graph representations under limited computational budgets.
Appendix A Methodologies
A.1. Dataset
@@
Table 8. Dataset Information. class. represents classification task and reg. represents regression task.
@@
Table 9. Dataset statistics of three categories from Amazon e-commerce networks
@@ A.2. Pre-training Metric
  • For the community detection task, we set $\epsilon$ as $1$. The predicted label between the $i$-th node and the $j$-th node $\hat{Y}_{i,j}^{c}$ is $1$ if $\operatorname{sim}(i,j)\geq 0.5$ and 0 otherwise. Accuracy is calculated by comparing the predicted label $\hat{Y}_{i,j}^{c}$ with the ground truth labels $Y_{i,j}^{c}$ and is defined as the proportion of correctly predicted labels out of all possible node pairs:
    (3) $\text{Accuracy(CD)}=\frac{\sum_{(i,j)\in V}\mathbbm{1}(\hat{Y}_{i,j}^{c}=Y_{i,j}^{c})}{|V|(|V|-1)/2},$
    where $\mathbbm{1}(\cdot)$ is an indicator function and |V|(|V|1)/2|V|(|V|-1)/2 is the total number of unique node pairs in the graph. This metric effectively measures how well the model can identify community structures by correctly classifying node pairs as being in the same community or in different communities.
@@
    (4) $\text{Accuracy(GCL)}=\frac{\sum_{i,j}\mathbbm{1}(\hat{Y}_{G_{i},G_{j}}=Y_{G_{i},G_{j}})}{N},$
    where $\hat{Y}_{G_{i},G_{j}}$ is the predicted label indicating whether graph $G_{i}$ and $G_{j}$ are from the same dataset and $Y_{G_{i},G_{j}}$ is the ground truth label. $\hat{Y}_{G_{i},G_{j}}$ is $1$ if $\operatorname{sim}(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}})\geq 0$ and 0 otherwise.NN is the total number of evaluated graph pairs.
@@
    (5) $\text{MSE(SPD)}=\frac{1}{|E|}\sum_{(i,j)\in E}(h_{\text{SPD}}(\mathbf{P}^{L}_{i}\|\mathbf{P}^{L}_{j})-\text{SPD}_{i,j})^{2}.$
@@
    (6) $\text{MAE(MC)}=\frac{1}{|V|}\sum_{i\in V}\|h_{\text{MC}}(\mathbf{P}_{i}^{L})-Y_{i}^{m}\|_{1},$
    where $Y_{i}^{m}$ is the pre-computed label for the $i$-th node.
A.3. Uncertainty-based Loss Weighing
The scale of the loss of different tasks can be different, causing the overall loss to be dominated by a certain task, and ultimately the loss of the other tasks cannot affect the learning process of the network-sharing layers. We use the uncertainty-based loss-weighing method (kendall2018multi) to automatically balance the four pre-training tasks and unify the different scales. Moreover, the uncertainty value reflects the contribution of each task towards the overall pre-training process. A higher uncertainty value indicates a lower contribution (kendall2018multi). Let $\sigma_{\tau}$ and $\mathcal{L}_{\tau}$ represent the task-specific uncertainty value for the task $\tau$. The overall pre-training loss is computed by:
(7) $\mathcal{L}=$ $\frac{1}{\sigma_{\text{SPD}}^{2}}\mathcal{L}_{\text{SPD}}+\frac{1}{\sigma_{\text{MC}}^{2}}\mathcal{L}_{\text{MC}}+\frac{1}{\sigma_{\text{CD}}^{2}}\mathcal{L}_{\text{CD}}+\frac{1}{\sigma_{\text{GCL}}^{2}}\mathcal{L}_{\text{GCL}}$
A.4. Pre-training Setting